a) total overtime production cost
b)total inventory holding cost for January through August
c)total stockout cost
d)total cost,excluding normal time labor cost for plan D
Homework: Chapter 13: Aggregate Planning-B Problem 136 800100
Answer #1

Total production possible = Regular + OT = 1600 + 20%*1600 = 1600 + 320 = 1920

Formulas:

OT production = Demand - Production - Ending inventory of previous month, if the value is negative then OT = 0, if positive the value can be maximum 320 (20%*1600)

Ending inventory = Regular production + OT production + Ending inventory of previous month - Demand. If the value is negative it means we are not able to meet demand and ending inventory is zero.

For Example in Jan = 1600+2000+0-1500 = 300

May Ending inventory = 1600+320+20 - 2300 = -360 (Means it will be zero)

Stockouts = Demand - (Regular production + OT production + Ending inventory of previous month). if the value is negative it means there are no stockouts and hence stockout = zero.

For Example in May = 2300 - (1600+320+20) = 2300 - 340 = 360

Please note - In july we have produced 300 OT units as required per demand instead of maximum 320 capacity.

Plan D Cost
Month Demand Production OT production Ending inventory Stockouts Month Inventory cost = Ending inventory * 20 OT cost = 50*OT units Stockout cost = Stockout units * 60
December 200 December
January 1,500 1600 0 300 0 January $                        6,000.00 0.0 0.0
February 1,700 1600 0 200 0 February $                        4,000.00 0.0 0.0
March 1,800 1600 0 0 0 March 0.0 0.0 0.0
April 1,900 1600 320 20 0 April $                            400.00 $       16,000.00 0.0
May 2,300 1600 320 0 360 May 0.0 $       16,000.00 $                        21,600.00
June 2,200 1600 320 0 280 June 0.0 $       16,000.00 $                        16,800.00
July 1,900 1600 300 0 0 July 0.0 $       15,000.00 0.0
August 1,400 1600 0 200 0 August $                        4,000.00 0.0 0.0
14,700 12,800 1,260 Total $                      14,400.00 $       63,000.00 $                        38,400.00
total overtime production cost $                      63,000.00
Total inventory $                      14,400.00
Total Stockout cost $                      38,400.00
total cost,excluding normal time labor cost for plan D $                   115,800.00